Grade 12th passMechanics

find vector C such that A + B + C = 0
when vector A has a magnitude of 4m and 60 degrees above the x axis (vector in quadrant one)
and vector B has a magnitude of 3m and 20 degrees below the x axis (vector in quadrant 4)
Thank you!

Dalton
9 Years agoGrade 12th pass

1 Answer

Nancy Thakur
9 Years ago
As we have A + B + C = 0when vector A has a magnitude of 4m and 60 degrees above the x axis A+B+C=04+3+C=0C=-7mTherfore the value of is-7